OpenStudy (anonymous):

Show that \[\frac{ 10 }{ v }+\frac{ 10 }{ u }=4\] and hence\[v=\frac{ 5u }{ (2u-5 }\]

OpenStudy (zzr0ck3r):

multiply the 1st term(top and bottom) by u and the second term (top and bottom) by v and you get \[\frac{ 10u+10v }{ u*v } = 4\] then multiply the left hand side and the right hand side by u*v 10u+10v = 4*u*v now solve for v. can you do the rest?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[10u+10v=4uv\] \[10v-4uv=-10u\] \[v(10-4u)=-10u\] \[v=\frac{-10u}{10-4u}=\frac{10u}{4u-10}=\frac{5u}{2u-5}\]

OpenStudy (anonymous):

hello hollywood!

OpenStudy (zzr0ck3r):

good now?